OpenStudy (anonymous):

multiplying and dividing Rational Expressions problems. : 4y+30/y^2-3y X y-3/2y+15 4y+30/y^2-3y X y-3/2y+15, 2y+15/y(1-3) X 1/1 2y+15/y "little lost" How do I get "2/y" for my answer

OpenStudy (rogue):

\[\frac {4y + 30}{y^2 - 3y} * \frac {y - 3}{2y + 15}\]

OpenStudy (rogue):

\[\frac {4y + 30}{y(y - 3)} * \frac {y - 3}{2y + 15}\]

OpenStudy (rogue):

\[\frac {4y + 30}{y} * \frac {1}{2y + 15} = \frac {4y + 30}{y(2y+15)}\]

OpenStudy (rogue):

\[\frac {4y + 30}{y(2y+15)} = \frac {2(2y+15)}{y(2y+15)} = \frac {2}{y}\]

OpenStudy (anonymous):

thanks , Rogue